In one city, 51.5% of adults are female, 8.2% of adults are left-handed, and 5.2% are left-handed females. For an adult selected at random from the let
F = event that person is female
L = event person is left handed.
Find P (F or L) Round approximations the three decimal places.
1
Expert's answer
2013-02-12T09:47:22-0500
The probability of at least one of the two joint events is the sum of the probabilities of these events without probability of their co-occurrence
P( F or L) = P(F) + P(L) - P (FL)
P(F) - probability of the event F
P(L) - probability of the event P
P(FL) - probability of event (P&F)
In our case
P(F) = 0.515 - probability what adult is female
P(L) = 0.082 - probability what adult is left-handed
P(FL) = 0.052 - probability what adult is& left-handed female
P (F or L) = 0.515+0.082-0.052=0.545=54.5%
Answer:& P (F or L) =& 54.5%